# The Cobb–Douglas Production Function and the Old Bowley’s Law

### v3

@@ 1 Introduction
In the following discussion, we will revisit the so-called “Bowley’s law” [5, 6, 39]. Paul Samuelson introduced the term “Bowley’s law” in 1964, within the sixth American edition of his renowned textbook Economics [39], to describe the constant wage share observed in economic data. The term was a tribute to Sir Arthur Bowley, an English economist, mathematician, and statistician credited with pioneering the systematic collection and statistical analysis of wage data in the UK. Bowley had hypothesized as early as 1920 that the wage share might remain constant, and later, with Josiah Stamp, provided evidence supporting this hypothesis by comparing UK wage shares in 1911 and 1924. In his 1937 book Wages and Income in the United Kingdom since $1860$ [6], Bowley definitively stated the constancy of the wage share, a discovery that challenged the views of classical economists such as Ricardo. They traditionally regarded the proportions of land, capital, and labor as intrinsically variable.
@@ 3 Bowley’s law and the Cobb–Douglas production function
$Y=AL^{\beta}K^{1-\beta},$
and obtained a labor exponent result of $\beta=0.75$, which was later verified by the National Bureau of Economic Research to be $0.741$. In their subsequent work in the 1940s, they allowed for variable exponents on $K$ and $L$, leading to estimates that closely matched improved measures of productivity developed during that period (see for more details). The constant $A$ (total productivity) was determined to be $A=1.01$. In the final analysis, after observing how the Cobb–Douglas function was derived from a data set [9], it becomes apparent that the process was rooted in data mining. As a result, we can succinctly summarize it as follows:
@@
Under the assumption that the production function $Y$ in (2.2) is in the form (3.1), it is easy to verify the validity of Bowley’s law. Indeed, substituting (3.1) in (2.4), it follows that in this case $s_{L}=\beta$, that is wage (labor) share is constant and equal to the output elasticity of labor $\beta$. Similarly, we get for capital share $s_{K}=\alpha$.
@@ 4 The Cobb–Douglas function derived
Based on the information presented in Section 3, it can be concluded that Cobb and Douglas established in [9] the legitimacy of the production function (3.1) through a combination of economic reasoning, statistical methods, and empirical data. Specifically, they demonstrated that by assigning the values $\beta=0.75$, $\alpha=0.25$, and $A=1.01$ to the respective parameters, the production function enjoying constant returns to scale yielded a precise fit to the data set under consideration.
@@ 5 Wage share as a differential invariant
Indeed, let $\big{(}\mathbb{R}^{3},\pi,\mathbb{R}^{2}\big{)}$ be a trivial bundle, such that $\pi=pr_{1}$ and $(K,L,Y)$ be adapted coordinates. Then the corresponding jet bundles are $\big{(}J^{1}\pi,\pi_{1},\mathbb{R}^{2}\big{)}$ and $\big{(}J^{1}\pi,\pi_{1,0},\mathbb{R}^{3}\big{)}$, where (see the commuting diagram below)
@@
The differential operator (5.4) represents the infinitesimal action of a one-parameter Lie group in a 5-dimensional space. According to the fundamental theorem on invariants of regular Lie group actions (see [29, Chapter 8]), we should expect to determine $5-1=4$ fundamental invariants. Indeed, solving the partial differential equation ${\rm Pr}^{1}(\mathbf{u}_{2})(F)=0$ using the method of characteristics, we find $F=F(I_{1},I_{2},I_{2},I_{4})$, where the fundamental invariants $I_{i}$, $i=1,\ldots,4$, are found to be
@@ 6 The logistic growth model as a generalization
that make it fit to the data with RSS=428.27. We note that the parameters $b_{1}$ (labor), $b_{2}$ (capital), $b_{3}$ (production) presented in Table 1 and the parameters $\alpha$, $\beta$ in (6.12) satisfy the orthogonality condition (4.7) with good accuracy, as expected. It must be noted at this point that the function (6.8) with $N_{Y}=120$ and $N_{L}=N_{K}=150$ under the condition (3.2) yielded a much less accurate fit to this data in [44] with ${\rm RSS}=4336.976$. Finally, we compute the corresponding RSS for the Cobb–Douglas function (3.1) determined by the exponential growth parameters (4.23) fitted to the same data from the period 1947-2016 provided by the Federal Reserve Bank of St. Louis (https://fred.stlouisfed.org), employing the FRED tool, to find that in this case RSS = 584.4616 which is less than the RSS for the function (6.8) determined by the parameters presented in Table 1.
@@
$Y=g(L,K)=\frac{aL^{p}K^{1-p}}{1+bL^{p}K^{1-p}}$ (6.13)
was recently introduced, employing a heuristic approach, see, for example, Anita et al. [2], Capasso et al. [7], Engbers et al. [12], La Torre et al. [25] and the relevant references therein for more details and applications. Note that the production function (6.13) is reducible to the Cobb–Douglas function (4.2) (i.e., when $b=0$). Compare the function (6.13) to the Holling disc equation in eco-dynamics [18]. Also, we observe that the new production function (6.8) is reducible to the production function (6.13) when $K$ and $L$ $\ll$ $N_{K}$ and $N_{L}$ respectively, $N_{L},N_{K}\approx 1$, $C=1$ in (6.8) and $a=N_{Y}$, $b=1$ in (6.13).

### v4

@@ 1 Introduction
In the following discussion, we will revisit the so-called “Bowley’s law” [5, 6, 39]. Paul Samuelson introduced the term “Bowley’s law” in 1964, within the sixth American edition of his renowned textbook Economics [39], to describe the constant wage share observed in economic data. The term was a tribute to Sir Arthur Bowley, an English economist, mathematician, and statistician credited with pioneering the systematic collection and statistical analysis of wage data in the UK. Bowley had hypothesized as early as 1920 that the wage share might remain constant, and later, with Josiah Stamp, provided evidence supporting this hypothesis by comparing UK wage shares in 1911 and 1924. In his 1937 book Wages and Income in the United Kingdom since 186018601860 [6], Bowley definitively stated the constancy of the wage share, a discovery that challenged the views of classical economists such as Ricardo. They traditionally regarded the proportions of land, capital, and labor as intrinsically variable.
@@ 3 Bowley’s law and the Cobb–Douglas production function
$Y=AL^{\alpha}K^{1-\alpha},$
and obtained a labor exponent result of $\alpha=0.75$, which was later verified by the National Bureau of Economic Research to be 0.7410.7410.741. In their subsequent work in the 1940s, they allowed for variable exponents on $K$ and $L$, leading to estimates that closely matched improved measures of productivity developed during that period (see for more details). The constant $A$ (total productivity) was determined to be $A=1.01$. In the final analysis, after observing how the Cobb–Douglas function was derived from a data set [9], it becomes apparent that the process was rooted in data mining. As a result, we can succinctly summarize it as follows:
@@
Under the assumption that the production function $Y$ in (2.2) is in the form (3.1), it is easy to verify the validity of Bowley’s law. Indeed, substituting (3.1) in (2.4), it follows that in this case $s_{L}=\alpha$, that is wage (labor) share is constant and equal to the output elasticity of labor $\beta$. Similarly, we get for capital share $s_{K}=\beta$.
@@ 4 The Cobb–Douglas function derived
Based on the information presented in Section 3, it can be concluded that Cobb and Douglas established in [9] the legitimacy of the production function (3.1) through a combination of economic reasoning, statistical methods, and empirical data. Specifically, they demonstrated that by assigning the values $\alpha=0.75$, $\beta=0.25$, and $A=1.01$ to the respective parameters, the production function enjoying constant returns to scale yielded a precise fit to the data set under consideration.
@@ 5 Wage share as a differential invariant
[…]
@@
The differential operator (5.4) represents the infinitesimal action of a one-parameter Lie group in a 5-dimensional space. According to the fundamental theorem on invariants of regular Lie group actions (see [29, Chapter 8]), we should expect to determine 51=45-1=45 - 1 = 4 fundamental invariants. Indeed, solving the partial differential equation ${\rm Pr}^{1}(\mathbf{u}_{2})(F)=0$ using the method of characteristics, we find $F=F(I_{1},I_{2},I_{2},I_{4})$, where the fundamental invariants $I_{i}$, $i=1,\ldots,4$, are found to be
@@ 6 The logistic growth model as a generalization
that make it fit to the data with ${\rm RSS}=428.27$. We note that the parameters $b_{1}$ (labor), $b_{2}$ (capital), $b_{3}$ (production) presented in Table 1 and the parameters $\alpha$, $\beta$ in (6.12) satisfy the orthogonality condition (4.7) with good accuracy, as expected. It must be noted at this point that the function (6.8) with $N_{Y}=120$ and $N_{L}=N_{K}=150$ under the condition (3.2) yielded a much less accurate fit to this data in [44] with ${\rm RSS}=4336.976$. Finally, we compute the corresponding RSS for the Cobb–Douglas function (3.1) determined by the exponential growth parameters (4.23) fitted to the same data from the period 1947–2016 provided by the Federal Reserve Bank of St. Louis (https://fred.stlouisfed.org), employing the FRED tool, to find that in this case RSS = 584.4616 which is less than the RSS for the function (6.8) determined by the parameters presented in Table 1.
@@
$Y=g(K,L)=\frac{aK^{p}L^{1-p}}{1+bK^{p}L^{-p}}$ (6.13)
was recently introduced, employing a heuristic approach, see, for example, Anita et al. [2], Capasso et al. [7], Engbers et al. [12], La Torre et al. [25] and the relevant references therein for more details and applications. Note that the production function (6.13) is reducible to the Cobb–Douglas function (4.2) (i.e., when $b=0$). Compare the function (6.13) to the Holling disc equation in eco-dynamics [18]. Also, we observe that the new production function (6.8) is reducible to the production function (6.13) when $K$ and $L$ll $N_{K}$ and $N_{L}$ respectively, $N_{L},N_{K}\approx 1$, $C=1$ in (6.8) and $a=N_{Y}$, $b=1$ in (6.13).